zarkam21:

what is done to 77,000 , 53,000 and 24,000 to get the answer 45.28?

Vocaloid:

@Hero @SmokeyBrown I have no idea what's going on in this problem

SmokeyBrown:

I'm not sure I understand the question either. I notice that 77,000 is 53,000 plus 42,000... Are we trying to get 45.28 as a result by combining the other three numbers in some way?

zarkam21:

yes exactly

Bob:

42k + 53k is 77k

Bob:

and what kind of question is this

Bob:

24k not 42k

SmokeyBrown:

Not sure if this is what you mean, but here is something I tried: (77,000 * 24,000)/53,000 = 24,868 24,868/24,000 = 1.4548 1.4528 - (24,000/24000) = 0.4528. So I got the digits, 4528 in a kind of convoluted way (is it cheating to reuse the numbers?) But, yeah, I can't get any further than that.

zarkam21:

nope not cheating

Bob:

who's cheating

zarkam21:

sorry guys for the late replies (s well as short), running a high fever

zarkam21:

oh em gee @SmokeyBrown that is exactly what I needed

zarkam21:

Okay so you did get 4528?

SmokeyBrown:

Heh, not quite. I got 0.4528, which is off from 45.28 by a factor of 100 and off from 4528 by a factor of 10000

zarkam21:

yes well with the overall calculations you multiply it by 100 so you were right

SmokeyBrown:

Is that so? Well, you've got the question in front of you, so I guess if that's allowed, hooray.

zarkam21:

okay now we have to do it with 79,000 , 57,000 , 22,000

zarkam21:

you set forward the process so I will do it, if you can check it i would appreciate it

SmokeyBrown:

Oh, I guess if you're allowed to reuse numbers, you could just do like (24,000) + itself 100 times, then divide by 24000 to get 100. Seems kind of like a cop-out though. Like, you could do the same for any of the numbers, say 4328 times, then divide that by 100. Maybe I'm still not fully understanding the question :/

zarkam21:

it got 0.3862

zarkam21:

wby

zarkam21:

okay then lol

SmokeyBrown:

Whoops, sorry, I was working on a writing assignment. Let me know if you need help with anything else!

Hero:

Where does this question come from? Is it a homework question?

zarkam21:

Nope, I was trying to figure a problem out. It wasn't homework just a lesson